Anna's car can be driven 44 miles per gallon of gas. If she buys gas for $3.06 a gallon, how much will she have to pay for gas on a 500 mile trip?
Answer by stanbon(75887)

Anna's car can be driven 44 miles per gallon of gas. If she buys gas for $3.06 a gallon, how much will she have to pay for gas on a 500 mile trip?
----------
gallons = distance/(miles per gallon)
------
g = 500/44 = 13.64 gallons
---------------------------
Cost = price*gallons = 3.06*13.64 = $41.73
------------
Cheers,
Stan H.
--------------